OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (anonymous):

A. x = 4, y = 11 B. x = -12, y = -5 C. x = -4, y = 3 D. x = 8, y = 3

OpenStudy (anonymous):

Since y = -2x + 19 and y = x + 7 , therefore , -2x+19=x+7 = -2x-x = 7-19 = -3x = -12 = x= -12/-3 = x = 4

Parth (parthkohli):

Now solve for y by adding 7 to x.

OpenStudy (anonymous):

??